OpenStudy (anonymous):

Please help.. Find the height of an object 5 seconds after it has been projected upward at a rate of 120 feet per second.

OpenStudy (anonymous):

Assuming we're on earth, we know gravity is a constant -32 ft/s^2 a(t)=a v(t)=v(o)+at x(t)=x(o)+v(o)t+(1/2)at^2 x(o)=0, we know v(o), and we know a=-32 ft/s^2 now just plug in t=5s and solve for x

OpenStudy (anonymous):

The equation my book give me is h=-16t^2+v(o)t... Is this the same as what you wrote or is it different? Aslo with that equation h=-16*5^2+120(0)5 How do I work out the 120(0)5?

OpenStudy (anonymous):

@dannym

OpenStudy (dan815):

you want to find when h , when t = 5

OpenStudy (dan815):

the V(o) there doesnt mean velocity * zero it means Vo initial velocity

OpenStudy (anonymous):

I've got that..

OpenStudy (anonymous):

Ok so the V(o)t would be the 120*5?

OpenStudy (dan815):

yes

OpenStudy (anonymous):

Ok thank you!

OpenStudy (dan815):

oh and is that equation given to you, are are you expected to find it yourself

OpenStudy (dan815):

or are*

OpenStudy (anonymous):

It was given to me in my book.

OpenStudy (anonymous):

lol hey guess what 1/2 * 32 is????

OpenStudy (dan815):

ok never mind

OpenStudy (dan815):

dannym was trying to show you from integration how to derive any equation for these kinda questions when you dont take into account the frictional air forces

OpenStudy (anonymous):

Thank you both! Much appreciated!

OpenStudy (dan815):

yea sure thing..